OpenStudy (anonymous):

The line through (1, -4) and parallel to 2x +3y = 4.

OpenStudy (kc_kennylau):

Should we find the slope of the line first by expressing the equation of the form of \(y=mx+c\)?

OpenStudy (anonymous):

I think so but if I plug the points (1,-4) in then I get 2-12=4

OpenStudy (kc_kennylau):

Oh if they're parallel we only have to change the constant (the y-intercept) !!!! :D

OpenStudy (kc_kennylau):

So we have \(2x+3y=c\)

OpenStudy (kc_kennylau):

So \(c\) is \(2-12=-10\) :D

OpenStudy (anonymous):

ohh lol that makes sense !(:

OpenStudy (kc_kennylau):

This video will help you on solving more questions of the same kind :D https://www.youtube.com/watch?v=9hryH94KFJA

OpenStudy (kc_kennylau):

So the answer is \(2x+3y=-10\)

OpenStudy (anonymous):

alright thank you so muchh

OpenStudy (kc_kennylau):

no problem :)